If a body have initial velomcity of 10 m/s and it attains a velocity of 20 m/s in 5 seconds then its acceleration is how much m/s square?
1
Expert's answer
2014-04-07T13:04:55-0400
We use the formula v=v0+at, where v0 is the initial velocity, a is acceleration, v is the final velocity, t is time. So, a=(v-v0)/t=(20-10)/5=10/5=2 [m/s^2]